Use the function value given to determine the value of sin(0). Answer in exact form. cos(∅)=6/7
Select one:
a. 7/√85
b. √13/6
c. 6/√85
d. √13/7

Answer by Alan3354(69443)

Use the function value given to determine the value of sin(0). Answer in exact form. cos(∅)=6/7
-----------
sin(0) = 0
---
If you mean of the same angle:
Remember sin^2 + cos^2 = 1 for all angles.
--> sin = sqrt(1 - cos^2) = sqrt(1 - 36/49)
= sqrt(13/49)
= sqrt(13)/7